# Topic-Specific Classifiers are Better Relevance Judges than Prompted LLMs

### v1

@@ 6 Practical Considerations for Using Judge Adapters
To drive adoption of our approach within the wider research community, enabling the comparison of models on test collections that were previously not possible due to the retrieval of too many unjudged documents, we distribute code to train new judge adapters, as well as our existing judge adapters for DL19, DL20, and Robust04. However, we do so with the caveat that these models are only ever to be used for the specific use case of assessing unjudged documents for the intended topic. For practitioners looking to train judge adapters on new test collections, several considerations emerge from our experiments:
@@
The choice of base ranking model impacts both adapter effectiveness and computational requirements. Mono-decoders consistently outperformed other architectures in our experiments, with a justifiable increase in parameter budget. LoRA adapted monoT5 models are be able to be trained and ran for inference on most consumer GPUs, or even CPUs, in contrast to LLM-as-a-judge methods. Domain alignment between the base model’s pretraining corpus and the target test collection improves effectiveness, though cross-domain transfer remains feasible with slightly more training data to compensate for domain shift.
@@
Our experiments suggest judge adapters are most advantageous when: (1) a sufficient pool of existing judgments exists, as evaluation on relevance labels predicted by judge adapters provide more robust estimations of system effectiveness than alternative approaches; (2) deterministic, reproducible, and/or explainable judgments are required, as judge adapters provide a discriminative approach to automatic relevance judgment; and (3) scalability is required and/or inference cost are a concern, as LLM-as-a-judge approaches require significantly more computational resources, particularly with few-shot prompting.
Limitations.
Judge adapters are designed specifically to extend existing judgment pools by approximating how a particular assessor would judge previously unseen documents for a specific topic. Thus, the assessment of similar topics or using judge adapters for distillation is invalid. The inclusion of the models we have presented in this paper within a system in any way apart from judging whether a document is relevant to a topic or not at evaluation time constitutes the equivalent to ‘training on the test set’.
7 Conclusion
@@
The assumption that unjudged documents are non-relevant can lead to unreliable and incomparable results, especially as retrieval model effectiveness becomes far better than systems that contributed to judgment pools. We address this through topic-specific judge adapters: LoRA-based finetuning of pretrained rankers that deliberately overfit to individual assessors’ relevance criteria for individual topics. Aligning the predictions of pre-trained ranking models with topic-specific judgments addresses key limitations of previous work, specificially LLM-as-a-judge, namely circularity, amplification of biases, and extreme resource usage. Instead, offer both a reliable and scalable approach to extend shallow pools to greater depth through automatic judgment.
@@
Our experiments across three test collections demonstrate that judge adapters trained on as few as 128 examples per topic consistently achieve Spearman’s $\rho>0.87$ correlation with ground truth system rankings, substantially outperforming both the non-relevance assumption baseline and LLM-as-a-judge approaches. Judge adapters exhibit higher agreement with ground truth labels (Krippendorff’s $\alpha=0.69$ vs. $0.15$ to $0.49$ for LLMs), and avoid circularity by differing from ranking models through topic-specific weight adaptation. The use of LoRA enables lightweight distribution and deterministic evaluation at negligible inference cost. These results establish judge adapters as a practical solution that preserves human judgment as the gold standard while extending test collection utility. More broadly, our findings reveal that topic-specific alignment, capturing how a particular assessor judges a particular topic, is both feasible with limited training data and more effective than seeking general-purpose relevance classifiers or relying on LLMs’ pretrained relevance priors only.
@@
The ability to accurately encode human rater preferences about texts has benefits beyond being able to judge the relevance of an unjudged document in an information retrieval system without the involvement of the original human judge. Firstly, it enables the assessment of relevance from many more perspectives. We are able to very precisely model a human rater with relatively few judgments from them. However, relevance is a single dimension for evaluating the effectiveness of information retrieval systems. In many search scenarios dimensions like trustworthiness and understandability are key to evaluating effectiveness. We plan to investigate if the approach we developed in this paper can be used to model dimensions beyond relevance. Another major open problem facing the information retrieval community at the moment is the reliable offline evaluation of RAG responses, as these are generated at query time and thus not able to be evaluated within the Cranfield paradigm. Disregarding the evaluation of attribution and focusing exclusively on the textual content, we plan to investigate the same technique presented in this paper to evaluate the quality of RAG responses according to different utility dimensions.

### v2

@@ 6. Practical Considerations for Using Judge Adapters
To drive adoption of our approach within the wider research community, enabling the comparison of models on test collections that were previously not possible due to the retrieval of too many unjudged documents, we distribute code to train new judge adapters, and our existing judge adapters for DL19, DL20, and Robust04. However, we do so with the caveat that these models are only ever to be used for the specific use case of assessing unjudged documents for the intended topic. For practitioners looking to train judge adapters on new test collections, several considerations emerge from our experiments.
@@
The choice of base ranking model impacts both adapter effectiveness and computational requirements. Mono-decoders consistently outperformed other architectures in our experiments, with a justifiable increase in parameter budget. LoRA-adapted monoT5 models can be trained and run for inference on most consumer GPUs, or even CPUs, in contrast to LLM-as-a-judge methods. Domain alignment between the base model’s pretraining corpus and the target test collection improves effectiveness, though cross-domain transfer remains feasible with slightly more training data to compensate for domain shift.
@@
(dietz2025principles) identify fourteen evaluation tropes that can compromise the validity of LLM-based evaluation. Judge adapters inherently avoid several of these failure modes:
  1. (i)
    Circularity avoidance: Unlike LLM-as-a-judge approaches, where the same model can serve as both ranker and evaluator (dietz2025principles, Tropes #1–2), judge adapters are topic-specific classifiers that cannot function as general-purpose rankers. Their deliberate overfitting to individual topics prevents dual use.
  2. (ii)
    Preservation of assessor perspective: LLM judges risk homogenizing relevance judgments and losing variety of opinion (dietz2025principles, Trope #4). Judge adapters instead encode the specific relevance criteria of individual human assessors, preserving the diversity inherent in human judgment.
  3. (iii)
    Reproducibility and stability: LLM-based evaluators suffer from model evolution, where behavior changes across versions without notice (dietz2025principles, Trope #7). Judge adapters produce deterministic, versioned outputs that remain stable over time, enabling reproducible evaluation.
  4. (iv)
    Robustness to self-preference bias: LLMs tend to favor text from their own model family (dietz2025principles, Trope #3). Judge adapters are not subject to this bias, and instead closely reproduce human judgment.
  5. (v)
    Computational efficiency: When scalability is required or inference costs are a concern, judge adapters offer substantial advantages, as LLM-as-a-judge approaches require significantly more computational resources, particularly with few-shot prompting. However, the cost of initially assembling a human-sourced label set has to be acknowledged as bottleneck.
Limitations and Validity Constraints.
Judge adapters are designed specifically to extend existing judgment pools by approximating how a particular assessor would judge previously unseen documents for a specific topic. This design imposes strict validity constraints that must be observed to maintain evaluation integrity:
  1. (i)
    No cross-topic transfer: The assessment of similar topics or using judge adapters trained on one topic for another is invalid, violating the topic-specific alignment they are intended for.
  2. (ii)
    No system integration: The inclusion of judge adapters within a retrieval system, whether for re-ranking, filtering, or any purpose other than post-hoc evaluation, constitutes training on the test set. This directly instantiates the Circularity trope (dietz2025principles, Trope #1) and invalidates any resulting evaluation.
  3. (iii)
    No distillation or knowledge transfer: Using judge adapter predictions to train or improve retrieval models would create the feedback loops characteristic of Goodhart-style Overfitting (dietz2025principles, Trope #10), where systems optimize for the evaluation signal rather than genuine relevance.
  4. (iv)
    Test set leak considerations: Unlike LLMs that may have been trained on publicly available test collections (dietz2025principles, Trope #8), the base model for adapters (monoT5 in our case) has documented training data. However, practitioners training adapters on new collections should verify that their base model was not exposed to the target test collection.
  5. (v)
    Comparability to original judgments: Expanding the judgment pool makes comparisons to systems that previously treated unjudged documents as non-relevant incompatible. Both systems must be evaluated on the new, expanded judgments.
To support responsible use, we recommend releasing judge adapters alongside documentation specifying their intended scope. We encourage the community to treat these models as evaluation infrastructure rather than as components for system development, thereby preserving the separation between optimization and assessment that valid IR evaluation requires.
When to Prefer Alternatives.
Judge adapters are not universally optimal. LLM-as-a-judge approaches may be preferable when no existing human judgments are available for the target topic, precluding adapter training, or rapid prototyping across many new topics outweighs the need for precise alignment with specific assessor behavior. Practitioners should apply the guardrails recommended by (dietz2025principles) to mitigate these associated validity risks.
7. Conclusion
The assumption that unjudged documents are non-relevant can lead to unreliable and incomparable results, especially as retrieval models improve beyond the systems that contributed to judgment pools. However, using LLM-as-a-judge approaches instead to extend the labeled pool yields poor agreement with human judgment.
@@ 7. Conclusion
We address this shortcoming through topic-specific judge adapters: LoRA-based fine-tuning of pretrained rankers that deliberately overfit to individual assessors’ relevance criteria for individual topics. Aligning the predictions of pre-trained ranking models with topic-specific judgments addresses key limitations of previous LLM-based approaches, namely circularity, amplification of biases, and extreme resource usage. Judge adapters offer a reliable and scalable way to extend shallow pools to greater depth through automatic judgment.
Our experiments across three test collections demonstrate that judge adapters trained on as few as 64–256 examples per topic achieve strong correlation with ground truth system rankings, reaching $\rho\geq 0.96$ for NDCG, Precision, and Recall at $k=50$, with more moderate but consistent advantages at $k=10$ ($\rho\geq 0.84$). Paired significance tests confirm that adapters significantly outperform all LLM-as-a-judge configurations at higher evaluation depths ($p<0.001$), while differences at $k=10$ are in favor of adapter approaches but not always statistically significant. Judge adapters further exhibit substantially higher agreement with ground truth labels (Krippendorff’s $\alpha=0.81$$0.88$ vs. 0.320.320.580.58 for LLMs). Notably, reasoning-capable LLMs do not outperform simpler models on this task, suggesting relevance judgment is a calibration task rather than a reasoning task. While LLMs maintain moderately high system ranking correlations (up to $\rho=0.90$ for the most effective configurations), they systematically over-label documents as relevant. Yet, they do so consistently across systems, preserving relative orderings.
@@
These results establish judge adapters as a practical solution that preserves human judgment as the gold standard while extending test collection utility. More broadly, our findings reveal that topic-specific alignment, i.e., capturing how a particular assessor judges a particular topic, is both feasible with limited training data and more effective than seeking general-purpose relevance classifiers or relying on LLMs’ pretrained relevance priors. Through topic-specific weight adaptation and the use of LoRA, topic adapters offer lightweight distribution and deterministic evaluation at negligible training and inference cost, while avoiding the evaluation tropes (dietz2025principles) that compromise LLM-as-a-judge approaches.
@@
The ability to accurately encode human rater preferences about texts has benefits beyond being able to judge the relevance of an unjudged document in an information retrieval system without the involvement of the original human judge. Firstly, it enables the assessment of relevance from many more perspectives. We are able to very precisely model a human rater with relatively few judgments from them. Given that in many search scenarios dimensions like trustworthiness and understandability are key to evaluating effectiveness, we plan to investigate if the approach we developed in this paper can be used to model dimensions beyond topical relevance. Another major open problem facing the information retrieval community at the moment is the reliable offline evaluation of RAG responses, as these are generated at query time and thus not able to be evaluated within the Cranfield paradigm. Disregarding the evaluation of attribution and focusing exclusively on the textual content, we plan to investigate the same technique presented in this paper to evaluate the quality of RAG responses according to different utility dimensions.